# SI-MIL: Taming Deep MIL for Self-Interpretability in Gigapixel Histopathology

### v1

@@ 4.2 Slide-level classification performance
Baselines using pathological features: These baselines denote training Additive ABMIL with pathological features. To induce interpretability, we train MIL with pathological features, referred as PathFeat. However, this framework is non-interpretable as the projector $H(\cdot)$ maps the pathological features into a non-interpretable deep feature space. Therefore, we include a true interpretable baseline by training the MIL without $H(\cdot)$.
2-stage training using pathological features: Here, we first train the Additive ABMIL and extract top-$K$ attended patches for each WSI. Then, a self-interpretable linear classifier using the pathological features from the patches is trained. Specifically, we train the SI branch independent of the MIL branch, i.e., without PAG Top-$K$. This is analogous to the post-hoc analytical methods in [15, 10].
@@
Results: As observed in Table 1, conventional MIL using pathological features, and particularly the one without projector, performs considerably worse than the methods using deep features. This accuracy-interpretability trade-off often undermines the benefits of using interpretable frameworks/features. SI-MIL aims to close this performance gap by utilizing deep feature-based guidance. We find that SI-MIL, despite imposing a linear constraint (Eq. 9) on predictions, elevates the performance of pathological features to be on par with deep feature-based baselines.
@@
Ablation studies: In Table 1, we mainly showcase the significance of the meticulously designed components of SI-MIL. We show the implication of the PAG Top-$K$ module by omitting the perturbed Top-$K$ selection and blocking gradient flow from the SI branch to the MIL branch. In Table 1, we observe that a non-differentiable approach degrades the performance. This indicates that the most discriminative region identified by the MIL is potentially less effective in the pathological feature space, thus highlighting the need to find regions discriminative in both feature spaces for enhancing the predictive power of SI branch. It can also be observed that in both settings, with or without perturbed Top-$K$, knowledge distillation is instrumental in enhancing the performance. $\mathcal{L}_{KD}$ acts as a regularizer for the SI branch, pushing it to stay as close as to the high-performing MIL branch. Additional ablations demonstrating the effect of varying $K$ in the PAG Top-$K$ module, the number of PF-Mixer layers, and the percentile and temperature for scaling $\beta$ are presented in the Supp. Sec. 9.
@@ 5.1 Local Interpretation: Slide-level
SI-MIL can explain model predictions at WSI-level without relying on post-hoc methods [4]. Contrary to existing MIL [36, 56, 24], SI-MIL can produce both patch- and feature-level explanations, due to the linear mapping between the pathological features and output predictions. Fig. 3 presents aggregated patch-feature importance reports generated by SI-MIL for two TCGA-BRCA WSIs, elucidating the rationale behind the predictions. Below, we explain the setup for generating such reports and then quantify their quality in terms of user-friendliness and faithfulness.
@@
User-friendliness: We qualitatively evaluate the utility of the patch-feature reports by an expert pathologist. First, we presented top $K$ ($=20$) patches and corresponding nuclear maps for the WSIs in Fig. 3a to the expert. The selected IDC and ILC patches demonstrated good agreement with class-specific prior knowledge. The IDC patches contained coherent cancer cells forming malignant glands, nests, or sheets with commentary about nuclear size, shape, color, and chromatin texture; and the ILC patches showed infiltrating small round cells in single file configurations. Afterwards, the top 10 contributing pathological features and associated feature distributions, as identified by SI-MIL in Fig. 3c, were evaluated by the expert to assess their correlation with domain knowledge in classifying IDC and ILC. 90% and 80% of these features for IDC and ILC WSIs in Fig. 3a were found relevant, respectively, due to their strong association with cell cohesiveness, nuclear hyperchromaticity, and morphology of cancerous nuclei properties. These analyses helped the pathologist in reasoning with the model’s rationale and developing trust in model’s predictions. Interestingly, the pathologist commented on the utility of such feature-level relevance report in downstream correlations with genomic and laboratory data.
Figure 4: Cohort-level Interpretation: Separability of top $K$ patches of WSIs across classes in the pathological feature space. Multivariate and Univariate analyses depict that the top $K$ patches selected by SI-MIL and their pathological features are more separable.
Faithfulness: We evaluated the faithfulness of our reports by quantifying the alignment of the top identified pathological features with pathologist’s assessments. The evaluation involved the pathologist assigning relevance scores to the top features. Specifically, we selected 10 WSIs each from IDC and ILC, and generated patch-feature reports including top 10 contributing pathological features. Then, the reports were analyzed and the features were categorized into high-, moderate-, or non-relevant categories by the expert. The mean and standard deviation of the number of features in each category are reported separately in Table 3. Also, an aggregated percentage of the number of features in each category is reported. The analysis shows that the majority of the identified features are either highly or moderately relevant towards correct classification and interpretability. Among the non-relevant features, a few are interesting to be analyzed on larger cohorts to potentially discover new diagnostic biomarkers. The selection of some of the non-relevant features may also be due to certain misclassifications by HoVer-Net. This is left for future exploration.
@@ 5.2 Global Interpretation: Cohort-level
In this section, we holistically analyze how SI-MIL interprets at a global cohort-level and benefits over conventional MIL. Specifically, we perform univariate and multivariate statistical analysis to measure class-separability in the pathological feature space, inline with  [15, 10, 43].
@@
Univariate and Multivariate class-separability: Through global cohort-level analysis, we demonstrate that SI-MIL, which includes the co-learning of MIL and SI branches, optimizes the selection of more informative patches than conventional MIL. During inference for both the models, we separately collect the top $K$ attended patches across WSIs corresponding to the two classes in TCGA-BRCA. Subsequently, we use the pre-extracted $d$ pathological features for the selected patches, as described in Sec. 3.3. Formally, given $N_{1}$ and $N_{2}$ number of WSIs in the two classes, we construct pathological feature matrices $F_{1}\in\mathbb{R}^{(N_{1}\times K)\times d}$ and $F_{2}\in\mathbb{R}^{(N_{2}\times K)\times d}$ for both the models.
@@
Multivariate analysis employs t-SNE [60] to project $F_{1}$ and $F_{2}$ into a 2D embedding space, as shown in Fig. 4a. Afterwards, we measure the class-separability in terms of two metrics: (1) JSdiv@i, which entails fitting a 2D Gaussian mixture model with $i$ components to each class and calculates the Jansen-Shannon (JS) divergence between the two distributions; and (2) Silhouette score [50], an unsupervised metric to evaluate the quality of class-wise created clusters. Both the measures are distance-based metrics that aim to highlight how separable the patches from the two classes are, in the projected embedding space. To account for modeling variability, we report the mean and standard deviation of the metrics across 5-fold cross-validation, as presented in Sec. 4.1. It can be observed in the table in Fig. 4a that SI-MIL consistently provides higher class-separability scores than conventional MIL method. This can be attributed to the co-learning technique in SI-MIL, which results in selecting more informative patches for individual classes that are better separable in the pathological feature space.
@@
Univariate analysis examines the class-separability of patches for individual pathological features. For a given feature, i.e., a column in $F_{1}$ and $F_{2}$, we create class-wise density distributions and measure the JS divergence. For visual simplicity, we show the univariate analysis for the two pathological features for both SI-MIL and the conventional MIL in Fig. 4b. We can observe that the class-wise density distributions in SI-MIL are significantly better separated than the MIL. This further supports our argument of better patch selection in SI-MIL from multivariate analysis. For an aggregated univariate analysis, we rank the features by the decreasing order of JS divergence, and plot the median JS divergence against the increasing number of features. Similar to multivariate analysis, we state the mean and std of the medians across 5-fold cross-validation (Fig. 4b). We can observe that SI-MIL provides significantly better median class-separability for a good number of features, which strongly supports the enhanced quality of selected patches while preserving pathological understanding.
@@ 5.3 Dataset contribution
We contribute a comprehensive dataset aimed at enhancing interpretability and reproducibility in MIL research. It comprises of nuclei maps, pathological features, and SI-MIL-generated patch-feature importance reports for 2.2K WSIs. WSI processing and feature extraction involved significant computing resources (details in Supp. Sec. 11). The complete list of pathological features, including cell shape and texture properties, spatial configurations, and interactions among different cell types, is detailed in the supplementary material. We provide the key elements to enable researchers to further expand on the already comprehensive set.
@@ 6 Conclusion
We present Self-Interpretable MIL, which not only augments model interpretability by identifying salient regions and providing feature-level contributions within these regions but also achieves high performance on gigapixel WSI tasks. SI-MIL bridges the gap between AI-driven analysis and pathologist-friendly reasoning, a first of its kind in histopathology. From an evolutionary perspective, different cancers may share fundamentally similar characteristics; the pathological features in SI-MIL can capture these properties, possibly lending itself well to rare/unseen cancers. Future work will also involve integration of LLM-driven pathological concepts in model training.
@@
  • Hand-crafted pathological feature extraction (15)
@@ 7 Dataset details (additional)
We benchmark our SI-MIL on three WSI datasets, namely TCGA-BRCA, TCGA-Lung, and TCGA-CRC. SI-MIL necessitates pathological features for interpretable prediction. We use HoVer-Net for segmenting and classifying the nuclei, and afterwards computed hand-crafted pathological features. Since HoVer-Net is trained exclusively on 40$\times$ magnification patches, our analysis is confined to WSIs having 40$\times$ magnification. This ensures accurate nuclei prediction and thereby meaningful pathological feature extraction.
@@ 8.3 Interpretability analysis setup
For interpretability analysis, we compare the separability of the top $K$ patches in the pathological feature space between the conventional MIL and SI-MIL (refer to Figure 4). To ensure a fair comparison, we select WSIs from the held-out test set where both MIL methods result in correct predictions.
Note that, we employ 5-fold cross-validation on the training split and held out the test set. We chose the best-performing fold for both local (visualization and pathologist relevancy score experiment) and global (visualization) interpretability analysis for the MIL methods. However, for multivariate class-separability scores (refer to Figure 4), we report the median and standard deviation from all 5-folds. Similarly, we report the median and standard deviation of Jensen-Shannon (JS) divergence across all 5-folds in Figure  4.
8.4 SI-MIL complexity analysis:
The mitigation of the trade-off between performance and interpretability by SI-MIL can be attributed to the choice of pathological features and the SI-MIL design choices, instead of merely an increase in the number of model parameters. It can be justified by comparing the size and performance of SI-MIL with the competing baselines. The number of model parameters in SI-MIL is 625K, while those in conventional MIL with DINO/IN ViT-S, CTransPath and ReTCCL are 345K, 985K, and 5.25M, respectively. Despite the differences in model sizes, SI-MIL results in comparable performance with respect to the competing baselines, as shown in Table 1.
@@ 9 SI-MIL ablation studies: hyperparameters
Effect of varying $K$ in the PAG Top-$K$ module: In Figure 5, we illustrate the impact of varying $K$ on SI-MIL performance. We observe that a larger value of $K$ leads to a significant drop in performance compared to the default $K=20$. This decrease may be attributed to an increase in irrelevant noisy patches, which makes it difficult for the model to classify WSIs in the pathological feature space.

### v2

@@ 4.2 Slide-level classification performance
Baselines using PathExpert features: These baselines denote training Additive ABMIL with PathExpert features. To induce interpretability, we train MIL with PathExpert features, referred as PathFeat. However, this framework is non-interpretable as the projector $H(\cdot)$ maps the PathExpert features into a non-interpretable deep feature space. Therefore, we include a true interpretable baseline by training the MIL without $H(\cdot)$.
2-stage training using PathExpert features: Here, we first train the Additive ABMIL and extract top-$K$ attended patches for each WSI. Then, a self-interpretable linear classifier using the PathExpert features from the patches is trained. Specifically, we train the SI branch independent of the MIL branch, i.e., without PAG Top-$K$. This is analogous to the post-hoc analytical methods in [15, 10].
@@
Results: As observed in Table 1, conventional MIL using PathExpert features, and particularly the one without projector, performs considerably worse than the methods using deep features. This accuracy-interpretability trade-off often undermines the benefits of using interpretable frameworks/features. SI-MIL aims to close this performance gap by utilizing deep feature-based guidance. We find that SI-MIL, despite imposing a linear constraint (Eq. 9) on predictions, elevates the performance of PathExpert features to be on par with deep feature-based baselines.
@@
Ablation studies: In Table 1, we mainly showcase the significance of the meticulously designed components of SI-MIL. We show the implication of the PAG Top-$K$ module by omitting the perturbed Top-$K$ selection and blocking gradient flow from the SI branch to the MIL branch. In Table 1, we observe that a non-differentiable approach degrades the performance. This indicates that the most discriminative region identified by the MIL is potentially less effective in the PathExpert feature space, thus highlighting the need to find regions discriminative in both feature spaces for enhancing the predictive power of SI branch. It can also be observed that in both settings, with or without perturbed Top-$K$, knowledge distillation is instrumental in enhancing the performance. $\mathcal{L}_{KD}$ acts as a regularizer for the SI branch, pushing it to stay as close as to the high-performing MIL branch. Additional ablations demonstrating the effect of varying $K$ in the PAG Top-$K$ module, the number of PF-Mixer layers, and the percentile and temperature for scaling $\beta$ are presented in the Supp. Sec. 11.
@@ 5.1 Local Interpretation: Slide-level
SI-MIL can explain model predictions at WSI-level without relying on post-hoc methods [4]. Contrary to existing MIL [36, 56, 24], SI-MIL can produce both patch- and feature-level explanations, due to the linear mapping between the PathExpert features and output predictions. Fig. 3 presents aggregated patch-feature importance reports generated by SI-MIL for two TCGA-BRCA WSIs, elucidating the rationale behind the predictions. Below, we explain the setup for generating such reports and then quantify their quality in terms of user-friendliness and faithfulness.
@@
User-friendliness: We qualitatively evaluate the utility of the patch-feature reports by an expert pathologist. First, we presented top $K$ ($=20$) patches and corresponding nuclear maps for the WSIs in Fig. 3a to the expert. The selected IDC and ILC patches demonstrated good agreement with class-specific prior knowledge. The IDC patches contained coherent cancer cells forming malignant glands, nests, or sheets with commentary about nuclear size, shape, color, and chromatin texture; and the ILC patches showed infiltrating small round cells in single file configurations. Afterwards, the top 10 contributing PathExpert features and associated feature distributions, as identified by SI-MIL in Fig. 3c, were evaluated by the expert to assess their correlation with domain knowledge in classifying IDC and ILC. 90% and 80% of these features for IDC and ILC WSIs in Fig. 3a were found relevant, respectively, due to their strong association with cell cohesiveness, nuclear hyperchromaticity, and morphology of cancerous nuclei properties. These analyses helped the pathologist in reasoning with the model’s rationale and developing trust in model’s predictions. Interestingly, the pathologist commented on the utility of such feature-level relevance report in downstream correlations with genomic and laboratory data.
Figure 4: Cohort-level Interpretation: Separability of top $K$ patches of WSIs across classes in the PathExpert feature space. Multivariate and Univariate analyses depict that the top $K$ patches selected by SI-MIL and their PathExpert features are more separable.
Faithfulness: We evaluated the faithfulness of our reports by quantifying the alignment of the top identified PathExpert features with pathologist’s assessments. The evaluation involved the pathologist assigning relevance scores to the top features. Specifically, we selected 10 WSIs each from IDC and ILC, and generated patch-feature reports including top 10 contributing PathExpert features. Then, the reports were analyzed and the features were categorized into high-, moderate-, or non-relevant categories by the expert. The mean and standard deviation of the number of features in each category are reported separately in Table 3. Also, an aggregated percentage of the number of features in each category is reported. The analysis shows that the majority of the identified features are either highly or moderately relevant towards correct classification and interpretability. Among the non-relevant features, a few are interesting to be analyzed on larger cohorts to potentially discover new diagnostic biomarkers. The selection of some of the non-relevant features may also be due to certain misclassifications by HoVer-Net. This is left for future exploration.
@@ 5.2 Global Interpretation: Cohort-level
In this section, we holistically analyze how SI-MIL interprets at a global cohort-level and benefits over conventional MIL. Specifically, we perform univariate and multivariate statistical analysis to measure class-separability in the PathExpert feature space, inline with  [15, 10, 43].
@@
Univariate and Multivariate class-separability: Through global cohort-level analysis, we demonstrate that SI-MIL, which includes the co-learning of MIL and SI branches, optimizes the selection of more informative patches than conventional MIL. During inference for both the models, we separately collect the top $K$ attended patches across WSIs corresponding to the two classes in TCGA-BRCA. Subsequently, we use the pre-extracted $d$ PathExpert features for the selected patches, as described in Sec. 3.2. Formally, given $N_{1}$ and $N_{2}$ number of WSIs in the two classes, we construct PathExpert feature matrices $F_{1}\in\mathbb{R}^{(N_{1}\times K)\times d}$ and $F_{2}\in\mathbb{R}^{(N_{2}\times K)\times d}$ for both the models.
@@
Multivariate analysis employs t-SNE [60] to project $F_{1}$ and $F_{2}$ into a 2D embedding space, as shown in Fig. 4a. Afterwards, we measure the class-separability in terms of two metrics: (1) JSdiv@i, which entails fitting a 2D Gaussian mixture model with $i$ components to each class and calculates the Jansen-Shannon (JS) divergence between the two distributions; and (2) Silhouette score [50], an unsupervised metric to evaluate the quality of class-wise created clusters. Both the measures are distance-based metrics that aim to highlight how separable the patches from the two classes are, in the projected embedding space. To account for modeling variability, we report the mean and standard deviation of the metrics across 5-fold cross-validation, as presented in Sec. 4.1. It can be observed in the table in Fig. 4a that SI-MIL consistently provides higher class-separability scores than conventional MIL method. This can be attributed to the co-learning technique in SI-MIL, which results in selecting more informative patches for individual classes that are better separable in the PathExpert feature space.
@@
Univariate analysis examines the class-separability of patches for individual PathExpert features. For a given feature, i.e., a column in $F_{1}$ and $F_{2}$, we create class-wise density distributions and measure the JS divergence. For visual simplicity, we show the univariate analysis for the two PathExpert features for both SI-MIL and the conventional MIL in Fig. 4b. We can observe that the class-wise density distributions in SI-MIL are significantly better separated than the MIL. This further supports our argument of better patch selection in SI-MIL from multivariate analysis. For an aggregated univariate analysis, we rank the features by the decreasing order of JS divergence, and plot the median JS divergence against the increasing number of features. Similar to multivariate analysis, we state the mean and std of the medians across 5-fold cross-validation (Fig. 4b). We can observe that SI-MIL provides significantly better median class-separability for a good number of features, which strongly supports the enhanced quality of selected patches while preserving pathological understanding.
@@ 5.3 Dataset contribution
We contribute a comprehensive dataset aimed at enhancing interpretability and reproducibility in MIL research. It comprises of nuclei maps, PathExpert features, and SI-MIL-generated patch-feature importance reports for 2.2K WSIs. WSI processing and feature extraction involved significant computing resources (details in Supp. Sec. 13). The complete list of PathExpert features, including cell shape and texture properties, spatial configurations, and interactions among different cell types, is detailed in the supplementary material. We provide the key elements to enable researchers to further expand on the already comprehensive set.
@@ 6 Conclusion
We present Self-Interpretable MIL, which not only augments model interpretability by identifying salient regions and providing feature-level contributions within these regions but also achieves high performance on gigapixel WSI tasks. SI-MIL bridges the gap between AI-driven analysis and pathologist-friendly reasoning, a first of its kind in histopathology. From an evolutionary perspective, different cancers may share fundamentally similar characteristics; the PathExpert features in SI-MIL can capture these properties, possibly lending itself well to rare/unseen cancers. Future work will also involve integration of LLM-driven pathological concepts in model training.
7 Acknowledgments
Reported research was supported by NIH 1R21CA258493-01A1, NSF IIS-2123920, IIS-2212046, and the Stony Brook Profund 2022 grant.
@@ 7 Acknowledgments
  • Hand-crafted PathExpert feature extraction (17)
@@ 8 Dataset details (additional)
We benchmark our SI-MIL on three WSI datasets, namely TCGA-BRCA, TCGA-Lung, and TCGA-CRC. SI-MIL necessitates PathExpert features for interpretable prediction. We use HoVer-Net for segmenting and classifying the nuclei, and afterwards computed hand-crafted PathExpert features. Since HoVer-Net is trained exclusively on 40$\times$ magnification patches, our analysis is confined to WSIs having 40$\times$ magnification. This ensures accurate nuclei prediction and thereby meaningful PathExpert feature extraction.
@@ 9.3 Interpretability analysis setup
For interpretability analysis, we compare the separability of the top $K$ patches in the PathExpert feature space between the conventional MIL and SI-MIL (refer to Figure 4). To ensure a fair comparison, we select WSIs from the held-out test set where both MIL methods result in correct predictions.
Note that, we employ 5-fold cross-validation on the training split and held out the test set. We chose the best-performing fold for both local (visualization and pathologist relevancy score experiment) and global (visualization) interpretability analysis for the MIL methods. However, for multivariate class-separability scores (refer to Figure 4), we report the median and standard deviation from all 5-folds. Similarly, we report the median and standard deviation of Jensen-Shannon (JS) divergence across all 5-folds in Figure 4.
9.4 SI-MIL complexity analysis
The mitigation of the trade-off between performance and interpretability by SI-MIL can be attributed to the choice of PathExpert features and the SI-MIL design choices, instead of merely an increase in the number of model parameters. It can be justified by comparing the size and performance of SI-MIL with the competing baselines. The number of model parameters in SI-MIL is 625K, while those in conventional MIL with DINO/IN ViT-S, CTransPath and ReTCCL are 345K, 985K, and 5.25M, respectively. Despite the differences in model sizes, SI-MIL results in comparable performance with respect to the competing baselines, as shown in Table 1.
10 SI-MIL additional results
Here we provide the mean and standard deviations for the main experiments (refer to Table 1) in Table 4.
Table 4: Results indicate the mean and standard deviation of 5-fold cross-validation on test set. All methods are trained with Additive ABMIL as base MIL. Int. denotes self-interpretability of a method.
Lung BRCA CRC
Int. Acc. AUC Acc. AUC Acc. AUC
IN ViT-S $0.859\pm 0.014$ $0.919\pm 0.004$ $0.929\pm 0.011$ $0.967\pm 0.005$ $0.891\pm 0.013$ $0.898\pm 0.018$
RetCCL $0.860\pm 0.008$ $0.935\pm 0.003$ $0.929\pm 0.011$ $\textbf{0.976}\pm\textbf{0.001}$ $0.889\pm 0.015$ $0.891\pm 0.047$
CTransPath $\textbf{0.904}\pm\textbf{0.003}$ $\textbf{0.967}\pm\textbf{0.002}$ $0.920\pm 0.023$ $0.974\pm 0.002$ $\textbf{0.906}\pm\textbf{0.010}$ $\textbf{0.897}\pm\textbf{0.023}$
DINO ViT-S $0.896\pm 0.003$ $0.957\pm 0.003$ $\textbf{0.937}\pm\textbf{0.012}$ $0.974\pm 0.005$ $0.904\pm 0.006$ $\textbf{0.897}\pm\textbf{0.014}$
PathFeat $0.830\pm 0.015$ $0.888\pm 0.009$ $0.885\pm 0.014$ $0.950\pm 0.005$ $\textbf{0.886}\pm\textbf{0.016}$ $0.818\pm 0.031$
PathFeat w/o $H(\cdot)$ $0.767\pm 0.018$ $0.837\pm 0.016$ $0.889\pm 0.012$ $0.914\pm 0.003$ $0.853\pm 0.013$ $0.720\pm 0.044$
2-stage training $0.865\pm 0.007$ $0.932\pm 0.009$ $0.908\pm 0.017$ $0.924\pm 0.019$ $0.876\pm 0.020$ $0.862\pm 0.036$
SI-MIL (ours) $\textbf{0.884}\pm\textbf{0.018}$ $\textbf{0.941}\pm\textbf{0.009}$ $\textbf{0.944}\pm\textbf{0.028}$ $\textbf{0.968}\pm\textbf{0.012}$ $0.884\pm 0.017$ $\textbf{0.910}\pm\textbf{0.018}$
Ablation study of SI-MIL components
w/o PAG Top-$K$ $0.859\pm 0.009$ $0.936\pm 0.011$ $0.915\pm 0.023$ $0.922\pm 0.026$ $0.876\pm 0.022$ $0.869\pm 0.024$
w/o KD $0.853\pm 0.010$ $0.915\pm 0.007$ $0.932\pm 0.016$ $0.951\pm 0.024$ $0.878\pm 0.024$ $0.830\pm 0.039$
w/o PAG Top-$K$ KD $0.857\pm 0.005$ $0.924\pm 0.005$ $0.915\pm 0.009$ $0.899\pm 0.013$ $0.879\pm 0.022$ $0.858\pm 0.036$
@@ 11 SI-MIL ablation studies: hyperparameters
Effect of varying $K$ in the PAG Top-$K$ module: In Figure 5, we illustrate the impact of varying $K$ on SI-MIL performance. We observe that a larger value of $K$ leads to a significant drop in performance compared to the default $K=20$. This decrease may be attributed to an increase in irrelevant noisy patches, which makes it difficult for the model to classify WSIs in the PathExpert feature space.